Solve by completing the square and match your results
3x squared + 5x -1 = 0

Answer by edjones(8007)

3x^2+5x-1=0
x^2+ 5x/3 - 1/3 =0
x^2+ 5x/3 =1/3
x^2+ 5x/3 + 25/36 = 25/36 + 1/3
x^2+ 5x/3 + 25/36 = 37/36
(x+ 5/6)^2=37/36
x+ 5/6 = +-sqrt(37)/6
x=-5/6 +-sqrt(37)/6
x=(-5+sqrt(37))/6, x=(-5-sqrt(37))/6
.
Ed